hiiiihiiiiiiiiii:

Subtract the fractions. Reduce the answers to lowest terms. 13/16 - 2/16 10/16 0 5/16 11/16 I thank its D

hiiiihiiiiiiiiii:

is it D

xxDeppressionxx:

1316216=? The two fractions have like denominators so you can subtract the numerators. Then: 13216=1116 This fraction cannot be reduced. Therefore: 1316216=1116 Solution by Formulas Apply the fractions formula for subtraction, to 1316216 and solve (13×16)(2×16)16×16 =20832256 =176256 Reduce by dividing both the numerator and denominator by the Greatest Common Factor GCF(176,256) = 16 176÷16256÷16=1116 Therefore: 1316216=11/16

hiiiihiiiiiiiiii:

so its D

xxDeppressionxx:

yes

MrMudd183:

i think so, just ask Deppression to do it for you.....i'm tired of giving my reasoning and getting it stolen before i even say it

hiiiihiiiiiiiiii:

ok ty

xxDeppressionxx:

np